OpenStudy (anonymous):

Newton binomial (2x - 1) ^4

OpenStudy (theviper):

\[(2x - 1) ^4\] \[(2x-1)^2(2x-1)^2\] \[(4x^2+1-4x)(4x^2+1-4x)\]

OpenStudy (theviper):

can u solve it now @RandyNemesis ???

OpenStudy (theviper):

@mayankdevnani right ???

OpenStudy (anonymous):

\[\left(\begin{matrix}4 \\ 1\end{matrix}\right)2x^{4}1^{0}-\left(\begin{matrix}4 \\ 2\end{matrix}\right)2x^{3}{1}^{1}\cdots\]

OpenStudy (anonymous):

The answer supousto look like this 32x^4-240x^3..... but nothouse numbers

OpenStudy (anonymous):

check your answers again. Or, if you don't believe the people on top, try this for a change. http://www.wolframalpha.com/input/?i=(2x-1)%5E4&t=crmtb01

OpenStudy (anonymous):

Thank you all! Shadowys u the best!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (mayankdevnani):

right @TheViper

OpenStudy (anonymous):